### 05 Mar 2024 v1

@@ Abstract
The three-body problem (3BP) poses a longstanding challenge in physics and celestial mechanics. Despite the impossibility of obtaining general analytical solutions, statistical theories have been developed based on the ergodic principle. This assumption is justified by chaos, which is expected to fully mix the accessible phase space of the 3BP. This study probes the presence of regular (i.e. non chaotic) trajectories within the 3BP and assesses their impact on statistical escape theories. Using numerical simulations, we establish criteria for identifying regular trajectories and analyse their impact on statistical outcomes. Our analysis reveals that regular trajectories occupy up to 32% of the phase space, and their outcomes defy the predictions of statistical escape theories. The coexistence of regular and chaotic regions at all scales is characterized by a multi-fractal behaviour. Integration errors manifest as numerical chaos, artificially enhancing the mixing of the phase space and affecting the reliability of individual simulations, yet preserving the statistical correctness of an ensemble of realizations. Our findings underscore the challenges in applying statistical escape theories to astrophysical problems, as they may bias results by excluding the outcome of regular trajectories. This is particularly important in the context of formation scenarios of gravitational wave mergers, where biased estimates of binary eccentricity can have significant consequences.
keywords:
@@ 1 Introduction
The gravitational three-body problem (3BP) serves as an exemplary case of chaos in nature, showcasing the complexity that can emerge in seemingly simple systems [1]. The inherently chaotic nature of the 3BP makes it challenging to find analytic solutions, with the only existing general solution being impractical because of slow convergence [2, 3, 4], and all other solutions being limited to periodic orbits [5, 6, 7, 8, 9, 10, 11, 12] or to the hierarchical configuration [13, 14].
Even obtaining individual solutions for specific realizations of the non-hierarchical 3BP is a far more complex task than it might initially appear. We define the non-hierarchical 3BP as the scenario in which three bodies of comparable masses (i.e. $m_{1}\sim m_{2}\sim m_{3}$, as opposed to the restricted 3BP where $m_{1}\ll m_{2},m_{3}$) engage in interactions within a confined region of space, exchanging energy and angular momentum (here we exclude braids and other periodic solutions). In the point-particle limit, these complex interactions continue until one particle is ejected, leaving behind a bound binary system [15, 16, 17].
No matter how complex a three-body interaction may appear, its evolution can be subdivided into two distinct states that the system alternates between (see Figure 1). The first state occurs when all three particles interact within a confined region of space, devoid of any fixed hierarchy among them. We refer to this state of the interaction as a ‘democratic resonance’, as the particles freely exchange energy and angular momentum among each other [see 18]. Democratic resonances cannot persist indefinitely, because this configuration is inherently unstable, and the system will eventually form a temporary hierarchy, transitioning into the second state, which we term ‘excursion’. Excursions occur when two bodies form a temporary binary and the third body recoils while remaining bound to the binary centre of mass. In this state, the system forms an unstable, temporary hierarchical triple, which will revert back to the democratic resonance state once the single body returns to interact with the binary. A three-body interaction cycles between these two states until the system finally breaks up into an unbound binary-single pair. The breakup can only occur from the democratic resonance state, because during an excursion, the system can be considered as non-interacting. The physical origin of chaos in the non-hierarchical 3BP arises from the close gravitational interactions during the democratic resonances.
@@
Fortunately, it is possible to leverage chaos to find statistical solutions of the 3BP, also called statistical escape theories, which have origin in the pioneering work of Monaghan [19]. His approach, which is based on the microcanonical ensemble of statistical mechanics, laid the foundation for an exploration of the interplay between deterministic chaos and probabilistic dynamics in celestial mechanics.
@@
The key idea of statistical escape theories is that chaotic interactions mix the phase space of the 3BP, so that over the lifetime of the system, every available state will be eventually visited, filling the phase space in a uniform and random sense. Except conserved quantities (energy and angular momentum vector), the system has lost memory of its initial conditions due to chaos and, at any given time, it can be found at any point in the accessible phase space. This is the definition of an ergodic system in statistical mechanics, meaning that the time average of some property over the trajectory is equal to the ensemble average over the phase space volume [20]. In the case of the non-hierarchical 3BP, it means that the probability of outcome properties are proportional to the phase space that is accessible at the breakup time of the triple system. Given energy and angular momentum vector, it is then possible to characterize and sample the accessible phase space at the moment of breakup to obtain the distributions of final outcome properties [19, 21].
@@
Statistical escape theories typically provide the closed-form expression for the tri-variate distribution of outcome properties $f(E_{\rm B},L_{\rm B},i_{\rm S})$, expressed in terms of final binary binding energy $E_{\rm B}$, angular momentum $L_{\rm B}$ and the angle $i_{\rm S}=\arccos(\vec{L}_{\rm S}\cdot\vec{L}_{\rm 0})$ between the total angular momentum $\vec{L}_{0}$ and the angular momentum of the ejected single $\vec{L}_{\rm S}$. The original statistical escape theory by Monaghan [19] has been improved and refined over the years, yielding a remarkably good agreement with the results of numerical experiments [22, 23, 24, 21, 25, 26, 27, 28, 29].
@@
However, statistical escape theories suffer from the assumption that the phase space of the 3BP is fully mixed. It has been highlighted that the 3BP is not fully chaotic, and its phase space is divided between chaotic and regular (i.e. non-chaotic) trajectories [30, 31, 32, 27, 33]. Therefore, we can expect that the predictions of the statistical escape theories clash with the outcomes of regular 3BP interactions. Only recently, a new statistical theory of the 3BP, not based on ergodicity, has been proposed with the aim of solving this issue [34]. Even though the theory is not complete yet, preliminary results show an improved agreement of the predicted ejection probabilities with the numerical experiments, compared to previous theories [28, 35].
@@
In this paper, we explore the presence of regular trajectories in the 3BP and its impact on the statistical predictions of the 3BP escape theories. Section 2 details the setup of our simulations and our numerical model. Criteria for identifying regular trajectories in the initial phase space are outlined in Section 3, which also illustrates the influence of regular regions on the statistical outcomes of the 3BP. Section 4 delves into the fine-grained mixing between regular and chaotic regions occurring at all scales, characterizing the multi-fractal nature of this mixing. Integration errors and their role in spurious phase space mixing, referred to as numerical chaos, are examined in Section 5. Section 6 highlights the extreme sensitivity of the 3BP to perturbations below the Planck length, where mixing between regularity and chaos persists. Finally, our conclusion and outlooks are presented in Section 7.
@@ 2 Numerical methods
We integrate the motion of the 3 particles assuming no particle size and no collision, by means of the tsunami integrator [36, 37]. tsunami uses a combination of algorithmic techniques [38, 39, 40] to improve the accuracy and speed of the $N$-body integration, which results in total energy errors $<10^{-10}$.
We also keep track of the state of the simulation using the same classification scheme outlined in Manwadkar et al. [35] and Trani et al. [41]. Namely, we stop the simulation when a single body is on a hyperbolic, diverging orbit with respect to the remaining binary, and we record this time as the lifetime $t_{\rm life}$ (alternatively, the disruption time) of the triple system. The simulations are integrated until either the system breaks up, or until we reach a total integration time of $10^{9}\rm\,yr$. Only 6 simulations out of $10^{6}$ have not reached the triple breakup by the final integration time.
@@ 3 Regular trajectories elude statistical escape theories
The top panel of Figure 3 shows the initial phase space map, color-coded by the identity of the escaping particle. In this space, neighbouring points share similar initial conditions. If the 3BP was fully ergodic, we should see a mix of colors everywhere, but instead we observe four large regions of uniform colors, two big ones at $\iota\sim 80^{\circ}$ and $260^{\circ}$, and two small ones at $\iota\sim 175^{\circ}$ and $\iota\sim 355^{\circ}$, that we term regular islands. These regular islands are constituted by regular trajectories (as opposed to chaotic trajectories), where the final state of the interaction smoothly maps onto the initial conditions. These correspond to escapes of either of the two members of the binary. The blue and green regular islands present a translational quasi-symmetry by ${}^{\circ}$ along the $\lambda$-axis, which correspond to the binary phase swapping the two binary members. The slightly unequal masses prevent it from being a perfect symmetry, and make the small green island (escape of the most massive $m_{1}=17.5$ particle) larger than the small blue island, and the big blue island more lopsided at high $\iota$ than the green one.
Figure 3: Initial condition space in inclination $\iota$ ($y$-axis) and true longitude $\lambda$ ($x$-axis). Each point represents an individual realization among $10^{6}$. Top panel: color-coded by the identity of the escaping particles. Red: $12.5\,\rm M_{\odot}$. Blue: $15\,\rm M_{\odot}$. Green: $17.5\,\rm M_{\odot}$. Bottom panel: fraction of same-color neighbouring particles $f_{\rm col}$, out of $k=12$ nearest neighbours.
To quantify how much area is occupied by the regular regions, we employ a $k$ nearest neighbour classification scheme. For each point in the phase space, we calculate the fraction $f_{\mathrm{col}}(k)$ of $k$ neighbours that have the same color as the point in consideration. The bottom panel of Figure 3 color-codes the realizations based on $f_{\rm col}$ for $k=12$. This criterion appears to effectively identify the major regular islands while also revealing numerous narrow, fine-grained regular regions outside these islands, which we refer to as regular stripes.
@@
As demonstrated by Manwadkar et al. [28], regular trajectories exhibit ejection probabilities that deviate from the predictions of statistical escape theories. For instance, if we consider all the simulations of Figure 3, the ejection probabilities for bodies of mass $12.5$, $15$, and $17.5\,\rm M_{\odot}$ are found to be 43%, 30%, and 27%, respectively. In contrast, according to the statistical escape theory of Stone and Leigh [26], these probabilities are expected to be 64%, 25%, and 11%, given our initial energy and angular momentum.
Figure 4: Distributions of semi-major axes (left panel) and eccentricity (right panel) of the final binary for the entire initial phase space (i.e. the one in Figure 3). The colors indicate the identity of the escaped particle, while the dashed lines indicate the predictions of the statistical escape theory of Stone and Leigh [26]. Thin lines: considering all simulations. Thick lines: excluding the regular islands ($f_{\rm col}<0.5$). In the left panel, we magnified the differences in semi-major axes distributions among different escaper by introducing different normalizations (×$\times 0.9$ for red, ×$\times 3$ for blue, ×$\times 20$ for green). In the right panel, the predictions for the eccentricity distributions for different escapers would be virtually identical, so we introduced different normalizations (×$\times 3$ for blue, ×$\times 5$ for green) to better distinguish the coloured lines.
An unexplored aspect thus far is how these regular trajectories influence the distribution of outcome properties. If we were to color-code the outcome properties such as the final binary semi-major axis in the initial phase space map, smooth transitions would appear corresponding to the regular regions. It is therefore not surprising that such regular regions have a considerable impact on the outcome properties, as shown in Figure 4, which compares the marginalized distributions of semi-major axis and eccentricity to the theoretical predictions of Stone and Leigh [26].
@@
Our findings have significant implications in the astrophysical application of statistical escape theories. Astrophysical predictions based on statistical theories will be inherently biased, because they exclude the outcomes of regular interactions. This issue is particularly concerning in the context of formation scenarios of gravitational wave mergers [e.g., 42, 43, 44], where the final binary eccentricity, known to dramatically affect the gravitational wave radiation mechanism (at high $e$, the coalescence time $t_{\rm GW}\propto(1-e^{2})^{7/2}$, see 45), is significantly different in the regular regions of the phase space.
@@ 4 Chasing Laplace’s demon
Similar patterns are visible if we color code by the lifetime of the system, as in Figure 6. Uniform regions in the ejected color map correspond to regions where the lifetime of the system is extremely short, comparable to the free fall time of the binary-single system $t_{\rm ff}=\frac{\pi}{\sqrt{G(m_{1}+m_{2}+m_{3})}}\left(\frac{d}{2}\right)^{3/2}\simeq 26\,\rm yr$. Regular interactions have very short lifetimes because the system breaks up after a single close encounter between all three-particles, which is not enough to make the system highly chaotic. Several studies have indeed found that there is an approximate power-law relation between the Lyapunov time $t_{\rm Lyap}$ and the lifetime $t_{\rm life}$ of the gravitational 3BP [46, 47, 48, 49]. The main reason is that a triple system with a longer lifetime has more opportunities to undergo democratic resonances than a short-lived one. On the other hand, a long-lived system may also have spent most of its time in an excursion state, so the relation between $t_{\rm Lyap}$ and $t_{\rm life}$ is only approximate [27].
@@
To better estimate the fractal nature of the phase space maps, we can measure the fractal dimension of every zoom-in. One of the most common ways to estimate the fractal dimension of a set of points is to calculate the box-counting dimension, or Minkowski–Bouligand dimension [50]. Unfortunately, this technique is not suited for our simulations, because we are limited by resolution at both the small scales (set by the average distance between realization in a box) and large scales (set by the outer boundary of each zoom-in box).
Therefore, we adopt instead the 2-point correlation dimension, which is known to be accurate even when only a small number of points is available [51]. This technique works by counting the number of pairs $C(l)$ whose distance is less then $l$. Then, for small $l$, $C(l)$ grows as a power-law, i.e. $C(l)\propto l^{D_{2}}$, where $D_{2}$ is the two-point correlation dimension, which is guaranteed to be close (strictly speaking, equal to or less than) the fractal dimension of the set.

### 08 Mar 2024 v2

@@ Abstract
Context:The three-body problem (3BP) poses a longstanding challenge in physics and celestial mechanics. Despite the impossibility of obtaining general analytical solutions, statistical theories have been developed based on the ergodic principle. This assumption is justified by chaos, which is expected to fully mix the accessible phase space of the 3BP.
Aims:This study probes the presence of regular (i.e. non chaotic) trajectories within the 3BP and assesses their impact on statistical escape theories.
Methods:Using numerical simulations, we establish criteria for identifying regular trajectories and analyse their impact on statistical outcomes.
Results:Our analysis reveals that regular trajectories occupy up to 32% of the phase space, and their outcomes defy the predictions of statistical escape theories. The coexistence of regular and chaotic regions at all scales is characterized by a multi-fractal behaviour. Integration errors manifest as numerical chaos, artificially enhancing the mixing of the phase space and affecting the reliability of individual simulations, yet preserving the statistical correctness of an ensemble of realizations.
Conclusions:Our findings underscore the challenges in applying statistical escape theories to astrophysical problems, as they may bias results by excluding the outcome of regular trajectories. This is particularly important in the context of formation scenarios of gravitational wave mergers, where biased estimates of binary eccentricity can significantly impact estimates of coalescence efficiency and detectable eccentricity.
Isles of regularity in a sea of chaos amid the gravitational three-body problem
Alessandro Alberto Trani 112233    Nathan W.C. ${}^{,}$ 4455    Tjarda C.N. ${}^{,}$ 66    Simon Portegies Zwart 77
(Received M DD, YYYY; accepted M DD, YYYY)
Key Words.:
@@ 1 Introduction
The gravitational three-body problem (3BP) serves as an exemplary case of chaos in nature, showcasing the complexity that can emerge in seemingly simple systems (Poincare, 1892). The inherently chaotic nature of the 3BP makes it challenging to find analytic solutions, with the only existing general solution being impractical because of slow convergence (Sundman, 1912; Belorizky, 1930; Barrow-Green, 2010), and all other solutions being limited to periodic orbits (Broucke & Boggs, 1975; Broucke, 1975; Hadjidemetriou & Christides, 1975; Hadjidemetriou, 1975; Henon, 1976; Moore, 1993; Chenciner & Montgomery, 2000; Šuvakov & Dmitrašinović, 2013) or to the hierarchical configuration (Kinoshita & Nakai, 1999, 2007).
Even obtaining individual solutions for specific realizations of the non-hierarchical 3BP is a far more complex task than it might initially appear. We define the non-hierarchical 3BP as the scenario in which three bodies of comparable masses (i.e. $m_{1}\sim m_{2}\sim m_{3}$, as opposed to the restricted 3BP where $m_{1}\ll m_{2},m_{3}$) engage in interactions within a confined region of space, exchanging energy and angular momentum (here we exclude braids and other periodic solutions). In the point-particle limit, these complex interactions continue until one particle is ejected, leaving behind a bound binary system (Agekyan & Anosova, 1971; Szebehely, 1971; Standish, 1972b).
No matter how complex a three-body interaction may appear, its evolution can be subdivided into two distinct states that the system alternates between (see Figure 1). The first state occurs when all three particles interact within a confined region of space, devoid of any fixed hierarchy among them. We refer to this state of the interaction as a ‘democratic resonance’, as the particles freely exchange energy and angular momentum among each other (see Heggie & Hut, 1993). Democratic resonances cannot persist indefinitely, because this configuration is inherently unstable, and the system will eventually form a temporary hierarchy, transitioning into the second state, which we term ‘excursion’. Excursions occur when two bodies form a temporary binary and the third body recoils while remaining bound to the binary centre of mass. In this state, the system forms an unstable, temporary hierarchical triple, which will revert back to the democratic resonance state once the single body returns to interact with the binary. A three-body interaction cycles between these two states until the system finally breaks up into an unbound binary-single pair. The breakup can only occur from the democratic resonance state, because during an excursion, the system can be considered as non-interacting. The physical origin of chaos in the non-hierarchical 3BP arises from the close gravitational interactions during the democratic resonances.
@@
Fortunately, it is possible to leverage chaos to find statistical solutions of the 3BP, also called statistical escape theories, which have origin in the pioneering work of Monaghan (1976a). His approach, which is based on the microcanonical ensemble of statistical mechanics, laid the foundation for an exploration of the interplay between deterministic chaos and probabilistic dynamics in celestial mechanics.
@@
The key idea of statistical escape theories is that chaotic interactions mix the phase space of the 3BP, so that over the lifetime of the system, every available state will be eventually visited, filling the phase space in a uniform and random sense. Except conserved quantities (energy and angular momentum vector), the system has lost memory of its initial conditions due to chaos and, at any given time, it can be found at any point in the accessible phase space. This is the definition of an ergodic system in statistical mechanics, meaning that the time average of some property over the trajectory is equal to the ensemble average over the phase space volume (Birkhoff, 1931). In the case of the non-hierarchical 3BP, it means that the probability of outcome properties are proportional to the phase space that is accessible at the breakup time of the triple system. Given energy and angular momentum vector, it is then possible to characterize and sample the accessible phase space at the moment of breakup to obtain the distributions of final outcome properties (Monaghan, 1976a, b).
@@
Statistical escape theories typically provide the closed-form expression for the tri-variate distribution of outcome properties $f(E_{\rm B},L_{\rm B},i_{\rm S})$, expressed in terms of final binary binding energy $E_{\rm B}$, angular momentum $L_{\rm B}$ and the angle $i_{\rm S}=\arccos(\@vec{L}_{\rm S}\cdot\@vec{L}_{\rm 0})$ between the total angular momentum ${L}_{0}$ and the angular momentum of the ejected single ${L}_{\rm S}$. The original statistical escape theory by Monaghan (1976a) has been improved and refined over the years, yielding a remarkably good agreement with the results of numerical experiments (Standish, 1972a; Saslaw et al., 1974; Valtonen, 1974; Monaghan, 1976b; Valtonen et al., 2005; Stone & Leigh, 2019; Manwadkar et al., 2020, 2021; Ginat & Perets, 2021).
@@
However, statistical escape theories suffer from the assumption that the phase space of the 3BP is fully mixed. It has been highlighted that the 3BP is not fully chaotic, and its phase space is divided between chaotic and regular (i.e. non-chaotic) trajectories (Shevchenko, 1998a, b, 2010; Manwadkar et al., 2020; Parischewsky et al., 2023). Therefore, we can expect that the predictions of the statistical escape theories clash with the outcomes of regular 3BP interactions. Only recently, a new statistical theory of the 3BP, not based on ergodicity, has been proposed with the aim of solving this issue (Kol, 2021). Even though the theory is not complete yet, preliminary results show an improved agreement of the predicted ejection probabilities with the numerical experiments, compared to previous theories (Manwadkar et al., 2021, 2023).
@@
In this paper, we explore the presence of regular trajectories in the 3BP and its impact on the statistical predictions of the 3BP escape theories. Section 2 details the setup of our simulations and our numerical model. Criteria for identifying regular trajectories in the initial phase space are outlined in Section 3, which also illustrates the influence of regular regions on the statistical outcomes of the 3BP. Section 4 delves into the fine-grained mixing between regular and chaotic regions occurring at all scales, characterizing the multi-fractal nature of this mixing. Integration errors and their role in spurious phase space mixing, referred to as numerical chaos, are examined in Section 5. Section 6 highlights the extreme sensitivity of the 3BP to perturbations below the Planck length, where mixing between regularity and chaos persists. In Section 7 we highlight the astrophysical implications of our work. Finally, our conclusion and outlooks are summarized in Section 8.
@@ 2 Numerical methods
We integrate the motion of the 3 particles assuming no particle size and no collision, by means of the tsunami integrator (Trani et al., 2019; Trani & Spera, 2023). tsunami uses a combination of algorithmic techniques (Stoer & Bulirsch, 1980; Mikkola & Aarseth, 1993; Mikkola & Tanikawa, 1999) to improve the accuracy and speed of the $N$-body integration, which results in total energy errors $<10^{-10}$.
We also keep track of the state of the simulation using the same classification scheme outlined in Manwadkar et al. (2023) and Trani et al. (2023). Namely, we stop the simulation when a single body is on a hyperbolic, diverging orbit with respect to the remaining binary, and we record this time as the lifetime $t_{\rm life}$ (alternatively, the disruption time) of the triple system. The simulations are integrated until either the system breaks up, or until we reach a total integration time of $10^{9}\rm\,yr$. Only 6 simulations out of $10^{6}$ have not reached the triple breakup by the final integration time.
@@ 3 Regular trajectories elude statistical escape theories
Figure 3: Initial condition space in inclination $\iota$ ($y$-axis) and true longitude $\lambda$ ($x$-axis). Each point represents an individual realization among $10^{6}$. Top panel: color-coded by the identity of the escaping particles. Red: $12.5\,\rm M_{\odot}$. Blue: $15\,\rm M_{\odot}$. Green: $17.5\,\rm M_{\odot}$. Regions of sparse density may appear as white points due to the white background. Bottom panel: fraction of same-color neighbouring particles $f_{\rm col}$, out of $k=12$ nearest neighbours.
The top panel of Figure 3 shows the initial phase space map, color-coded by the identity of the escaping particle. In this space, neighbouring points share similar initial conditions. If the 3BP was fully ergodic, we should see a mix of colors everywhere, but instead we observe four large regions of uniform colors, two big ones at $\iota\sim 80^{\circ}$ and $260^{\circ}$, and two small ones at $\iota\sim 175^{\circ}$ and $\iota\sim 355^{\circ}$, that we term regular islands. These regular islands are constituted by regular trajectories (as opposed to chaotic trajectories), where the final state of the interaction smoothly maps onto the initial conditions. These correspond to escapes of either of the two members of the binary. The blue and green regular islands present a translational quasi-symmetry by ${}^{\circ}$ along the $\lambda$-axis, which correspond to the binary phase swapping the two binary members. The slightly unequal masses prevent it from being a perfect symmetry, and make the small green island (escape of the most massive $m_{1}=17.5\,{{\rm M}_{\odot}}$ particle) larger than the small blue island, and the big blue island more lopsided at high $\iota$ than the green one.
Figure 4: Distributions of semi-major axes (left panel) and eccentricity (right panel) of the final binary for the entire initial phase space (i.e. the one in Figure 3). The colors indicate the identity of the escaped particle, while the dashed lines indicate the predictions of the statistical escape theory of Stone & Leigh (2019). Thin lines: considering all simulations. Thick lines: excluding the regular islands ($f_{\rm col}<2/3$). In the left panel, we magnified the differences in semi-major axes distributions among different escaper by introducing different normalizations (×$\times 0.9$ for red, ×$\times 3$ for blue, ×$\times 20$ for green). In the right panel, the predictions for the eccentricity distributions for different escapers would be virtually identical, so we introduced different normalizations (×$\times 3$ for blue, ×$\times 5$ for green) to better distinguish the coloured lines.
To quantify how much area is occupied by the regular regions, we employ a $k$ nearest neighbour classification scheme. For each point in the phase space, we calculate the fraction $f_{\mathrm{col}}(k)$ of $k$ neighbours that have the same color as the point in consideration. The bottom panel of Figure 3 color-codes the realizations based on $f_{\rm col}$ for $k=12$.ly This criterion appears to effectively identify the major regular islands while also revealing numerous narrow, fine-grained regular regions outside these islands, which we refer to as regular stripes.
@@
As demonstrated by Manwadkar et al. (2021), regular trajectories exhibit ejection probabilities that deviate from the predictions of statistical escape theories. For instance, if we consider all the simulations of Figure 3, the ejection probabilities for bodies of mass $12.5$, $15$, and $17.5\,\rm M_{\odot}$ are found to be 43%, 30%, and 27%, respectively. In contrast, according to the statistical escape theory of Stone & Leigh (2019), these probabilities are expected to be 64%, 25%, and 11%, given our initial energy and angular momentum.
An unexplored aspect thus far is how these regular trajectories influence the distribution of outcome properties. If we were to color-code the outcome properties such as the final binary semi-major axis in the initial phase space map, smooth transitions would appear corresponding to the regular regions. It is therefore not surprising that such regular regions have a considerable impact on the outcome properties, as shown in Figure 4, which compares the marginalized distributions of semi-major axis and eccentricity to the theoretical predictions of Stone & Leigh (2019).
@@
Our findings have significant implications in the astrophysical application of statistical escape theories. Astrophysical predictions based on statistical theories will be inherently biased, because they exclude the outcomes of regular interactions. This issue is particularly concerning in the context of formation scenarios of gravitational wave mergers (e.g., Kritos et al., 2022; Mapelli et al., 2022; Ginat & Perets, 2023), where the final binary eccentricity, known to dramatically affect the gravitational wave radiation mechanism (at high $e$, the coalescence time $t_{\rm GW}\propto(1-e^{2})^{7/2}$, see Peters, 1964), is significantly different in the regular regions of the phase space. We delve deeper into this aspect in Section 7.
@@ 4 Chasing Laplace’s demon
Similar patterns are visible if we color code by the lifetime of the system, as in Figure 6. Uniform regions in the ejected color map correspond to regions where the lifetime of the system is extremely short, comparable to the free fall time of the binary-single system $t_{\rm ff}=\frac{\pi}{\sqrt{G(m_{1}+m_{2}+m_{3})}}\left(\frac{d}{2}\right)^{3/2}\simeq 26\,\rm yr$. Regular interactions have very short lifetimes because the system breaks up after a single close encounter between all three-particles, which is not enough to make the system highly chaotic. Several studies have indeed found that there is an approximate power-law relation between the Lyapunov time $t_{\rm Lyap}$ and the lifetime $t_{\rm life}$ of the gravitational 3BP (Lecar et al., 1992; Orlov et al., 2010; Mikkola & Tanikawa, 2007; Urminsky & Heggie, 2009). The main reason is that a triple system with a longer lifetime has more opportunities to undergo democratic resonances than a short-lived one. On the other hand, a long-lived system may also have spent most of its time in an excursion state, so the relation between $t_{\rm Lyap}$ and $t_{\rm life}$ is only approximate (Manwadkar et al., 2020).
@@
To better estimate the fractal nature of the phase space maps, we can measure the fractal dimension of every zoom-in. One of the most common ways to estimate the fractal dimension of a set of points is to calculate the box-counting dimension, or Minkowski–Bouligand dimension (Schroeder & Wiesenfeld, 1991). Unfortunately, this technique is not suited for our simulations, because we are limited by resolution at both the small scales (set by the average distance between realization in a box) and large scales (set by the outer boundary of each zoom-in box).
Therefore, we adopt instead the 2-point correlation dimension, which is known to be accurate even when only a small number of points is available (Grassberger & Procaccia, 1983). This technique works by counting the number of pairs $C(l)$ whose distance is less then $l$. Then, for small $l$, $C(l)$ grows as a power-law, i.e. $C(l)\propto l^{D_{2}}$, where $D_{2}$ is the two-point correlation dimension, which is guaranteed to be close (strictly speaking, equal to or less than) the fractal dimension of the set.